Mr. Fuller has paid $144 in interest on a loan of 1,800 borrowed for a year. What is the rate of interest on the loan?
1
Expert's answer
2011-11-22T08:16:04-0500
Let x % be rate of interest (in percents). As Mr. Fuller has paid $144,
then:
144 = 1800*x/100.
So, we get x = 144*100/1800 = 8 %.